A simple rand om sample of 8 employees of a corporation provided the following information. Employee Age Gender 25 32 26 40 5024 22 23 M M M M F Determine a point estimate for the proportion of all employees who are female. .25 .50 0.375 None of the above
Answer #1

Here we have data:

n = 1+ 2 + 3 + 4 + 5 + 6 + 7 + 8

= 36

x = 5 + 6 + 8 = 19

Point estimation of the proportion of all employ who are female:

p = 19 / 36 = 0.53

Correct option " D"

None of the above